OpenStudy (anonymous):

A complex number, (a + bi), multiplied by (2 + 3i) and added to -i gives the product of (-11 + 5i) and (1 – i).

OpenStudy (anonymous):

A=? b-?

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@mathhasproblems

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@jabez177 @Jadeishere

OpenStudy (jadeishere):

I don't remember how to do this, I'm sorry, if I did I would be more than willing to help you

OpenStudy (anonymous):

Thank you.

OpenStudy (trojanpoem):

(a + bi) * (2 + 3i) -i = (-11 + 5i) * (1 – i) 2a + 3 ai + 2bi + 3bi^2 - i = -11 + 11i + 5i - 5 i ^2

OpenStudy (trojanpoem):

now we are doing to do sums

OpenStudy (trojanpoem):

then take all the imaginary numbers together and the others together.

OpenStudy (trojanpoem):

2a - 3b = -11 +5 3 ai + 2 bi - i = 11i + 5i

OpenStudy (trojanpoem):

Review after me.

OpenStudy (trojanpoem):

2 a - 3 b = -6 3 a + 2b = 15

OpenStudy (trojanpoem):

2a - 3 b = -6 * 2 3 a + 2b = 15 *-3

OpenStudy (trojanpoem):

4a - 6b = -12 -9a + 6b = -45

OpenStudy (trojanpoem):

I think i miscalculated something

OpenStudy (trojanpoem):

or the number isn't cool at all.

OpenStudy (anonymous):

Haha

OpenStudy (trojanpoem):

That's why I hate PCs. Give me a minute gonna do it on a paper.

OpenStudy (trojanpoem):

a = 3 b = 4

OpenStudy (trojanpoem):

2 a - 3 b = -6 the first EQN 3 a + 2 b = 17 the second EQN